OpenStudy (anonymous):

CAN ANYONE HELP ME WITH 2^6X=4,096

OpenStudy (anonymous):

plug in(replace X with) numbers till u get 4,096

OpenStudy (anonymous):

....

OpenStudy (rsadhvika):

hint - 4096 = 2^12

OpenStudy (anonymous):

x=4096/64=64

OpenStudy (anonymous):

thank you @rsadhvika and @eliassaab

OpenStudy (anonymous):

yw